Sunday 08/02/15

11:00 am – 3:30 pm, 6:30 pm – 11:15 pm


  • STORM2 froze during the night (click error provided an out of frame crash in python — built in check doesn’t catch it.
  • debugged this with Bogdan, changed catch statement to check size of the object queried, not just the expected size of the image.
  • restarted imaging of Ph-KD L2F03 with fresh buffer
  • acquisition still looks good at movie 18. And rolling.
  • need to take z-beads and chromatin beads tomorrow before Colenso starts.


  • need to design and order new Green probes.
  • need to analyze existing green data.


  • figuring out parsing of new data schema
  • trying to process L12 two-color data for analysis with Pipeline 5 or updated pipeline 1.

Nuclei live imaging

  • took time lapse images of Hoechst stained nulcei (2 year+ expired DRAQ5 didn’t look that good at 1:10,000)
  • took images of mock fix (replaced growth media with more growth media)
  • took images of formaldhyde fixed cells
  • took images of MeOH fixed cells (clearly shrunk)
  • need to do some coding to analyze these.
    • probably worth some manual playing with the Fiji tools too.
  • some issues with Z-focus and x-y registration, largely sorted though.


  • check that numbers don’t change much upon down-sampling


  • reading article, drafting comments of first impressions
  • very long methods section, reading imaging methods in detail
This entry was posted in Summaries. Bookmark the permalink.